Find fx(x,y), fy(x,y), fx(1,3), and fy(-2,4) for the given function. If 

𝑧 = 𝑓(𝑥, 𝑦) = 3𝑥ଷ𝑦ଶ − 𝑥ଶ𝑦ଷ + 4𝑥 + 9



1
Expert's answer
2021-06-09T17:10:27-0400

"z=f(x,y)=3xy-xy+4x+9"

"fx(x,y)=3y-y+4"

"fy(x,y)=3x-x"

"fx(1,3)=3(3)-3+4"

"\\=6+4\\=10"

"fy(-2,4)=3(-2)-(-2)\\=-6+2\\-4"